Aerospace

What is the maximum amount of quantization error that can occur in a 10bit system in the input analog voltage is from 0V to 6V?

a) 10mV

b) 6mV

c) 0.5mV

d) 1mV
Answer #1

Answer: b

Explanation: The minimum voltage step input is 6/(2N-1) = 6/1023 =5.865mV ≈6mV. This is the maximum error that can occur during the conversion process.


answered by: alibaser